A box with a square base and an open top is to have volume 62.5 m3. Neglect the thickness of the material used to make the box and find the dimension that will minimize the amount of material used.


1
Expert's answer
2022-02-07T16:54:03-0500

Let "x" be the length of the sides of the base and "h" be the height. Then,

We need to maximize;

"\displaystyle
SA=\text{area of base+4(area of vertical side)}=x^2+4xh"

subject to;

"V=\text{(area of base)(height)}=x^2h"

Now, "\displaystyle V=\text{(area of base)(height)}=x^2h\Rightarrow 62.5=x^2h\Rightarrow h=\frac{62.5}{x^2}"

"\displaystyle
\Rightarrow SA=x^2+4xh=x^2+4x\left(\frac{62.5}{x^2}\right)=x^2+250x^{-1}"

Thus, taking the derivative of "\displaystyle
SA=x^2+250x^{-1}" wrt "x" yields;

"(SA)\prime=2x-250x^{-2}"

Setting "(SA)\prime=0" yields;

"2x-250x^{-2}=0\Rightarrow x=5\ m"

The second derivative test verifies that "SA" has a minimum at "x=5\ m" since:

"\displaystyle
(SA)\prime\prime=2+500x^{-3}=6" which is positive at "x=5\ m"

Thus, at "x=5,\ SA=5^2+\frac{250}{5}=75"


Hence, the minimum surface area;

"SA=75\ m^2", and

the box should have base "5\ m" by "5\ m" and height "2.5\ m".